# Direct ab initio calculation of magnons in altermagnets: method, spin-space symmetry aspects, and application to MnTe

### v1

@@ Abstract
We suggest the method for direct ab initio calculation of magnons in complex collinear magnets. The method is based on the density-functional-theory (DFT) calculation under two different constraints: one constraint governs the change of the magnetization with respect to the ground state, and the other is the symmetry constraint responsible for the value of the magnon wave vector. The performance of the method is demonstrated by the application to an altermagnet MnTe. An important role in both the formulation and the application of the method play the aspects of generalized symmetry described by the spin-space groups. The symmetry analysis connects in one coherent picture the following three parts of the consideration: (i) the generalized translational symmetry of the magnons as a crucial condition for their efficient ab-initio calculation, (ii) altermagnetic spin-splitting of the electron states in the ground magnetic state, and (iii) chirality splitting of the magnon excitations. It is demonstrated that both the spin splitting of the electron states and the chirality splitting of the magnons, though very different in their physical nature, have identical patterns in the corresponding wave vector spaces. Since the altermagnetism of MnTe is the consequence of the presence of the Te atoms, an adequate attention is devoted to the symmetry analysis and calculation results for the Te moments induced in the magnon states. The knowledge on the symmetry properties of the Te moments allows to accelerate the numerical convergence of the magnon states and serves as a test for the accuracy of the calculations. To expose the connection between electron band structures of the magnon states of the system and the chirality properties of these states we investigate the transformation of the electron structure in the transition from the collinear ground state to a noncollinear magnon state.
@@ I Introduction
There are different approaches to the DFT based calculation of magnons. One of the approaches is the calculation of the dynamic magnetic susceptibility in nonuniform transversal magnetic field Savrasov1998 ; Buczek2011 ; Sandratskii2012 ; Odashima2013 ; Lounis2015 ; Skovhus2021 ; Liu2023 . A strong feature of this method is that it allows to investigate not only the structure and the energy of magnons but also their life time resulting from the interaction of magnons with single-electron Stoner excitations. In the case of complex magnetic materials the method is very demanding with respect to both computation techniques and computer resources. A widely used approach to the theoretical study of the adiabatic magnons is the mapping, as an intermediate step, of the electron system on the Heisenberg Hamiltonian of interacting atomic moments. Such a mapping is currently a standard procedure consisting in the DFT based evaluation of the Heisenberg exchange parameters (see, e.g., Refs. force_theorem, ; Mankovsky_parameters2022, ; Turek_parameters2006, ; Lezaic2006, ; Mu2019, ; Martinez2023, ; Sandratskii2007, ; Sodequist2024, ). An efficient method of the mapping was suggested by Liechtenstein et al. force_theorem and is based on the evaluation of the variation of the band energy as the response to the infinitesimal deviation of the atomic moments from the magnetization axis. The possibility to replace the variation of the total energy by the variation of the band energy is based on so-called magnetic force theorem force_theorem . A recent review of the development and applications of the Liechtenstein et al. method is given in Ref. Szilva2023, . In this paper we suggest a direct DFT-based method for the magnon calculation that does not include mapping of the electron system on the Heisenberg Hamiltonian.
The direct method needs more computer time than the mapping procedure by Liechtenstein et al. since each magnon state is calculated in a separate self-consistent DFT calculation. However, the method has features not provided by the standard mapping approach. Among them are the following: First, the fully self-consistent calculation of the magnon structure and energy is performed in contrast to the mapping approach based on the evaluation of the band energy variation from non self-consistent calculation. Second, in the magnon states of complex structures the atoms that are equivalent in the ground state may become inequivalent. The suggested method takes this into account in a consequent self-consistent manner. Third, in the compounds the contribution of the nominally nonmagnetic atoms to the magnon states is self-consistently taken into account. Forth, the method allows to estimate the dependence of the magnon energy on the number of magnons by varying the tilting angles of the atomic moments from the magnetization axis. The approach suggested by Liechtenstein et al. considers an infinitesimal deviation of the moments from the magnetization axis.
@@
derived in Ref. Niu99, for the magnon energy in an arbitrary collinear magnet. Here $\Delta E$ is the increase of the energy of the magnon state with respect to the GS and $\Delta m_{z}$ is the magnetization change with respect to the GS magnetization. Equation (1) correlates with the property that one magnon changes the magnetization of the system by 1 $\mu_{B}$. The first constraint imposes the condition on the magnetization of the system by means of introduction of an effective magnetic field. The second constraint specifies the value of the magnon wave vector. This is a symmetry type of constraint Sandratskii2001 where the invariance of the initial Hamiltonian with respect to the symmetry operation responsible for the desired property reproduces itself during iterations. The symmetry constraint does not need a constraining field. It is important to emphasize that to constrain the magnon wave vector the generalized translational periodicity described by the machinery of spin space groups (SSG) must be imposed.
@@
In the paper, we formulate the method in the form valid for the study of the magnons in any collinear magnet with weak spin-orbit coupling (SOC). The performance of the method is demonstrated by the application to altermagnets focusing on MnTe as a representative of this class of materials. As the notion of an altermagnet is rather new it is worth to introduce it briefly. The ground state of a two sublattice AFM is characterized by two mutually compensating ferromagnetic sublattices. In terms of the electron structure, the zero net magnetization of an AFM is the result of the mutual compensation of the spin-up and spin-down electron states which is the consequence of the symmetry properties of the material. Importantly, the realization of this spin compensation can be different for different AFM materials: It can take place either at each wave vector $\bf{k}$ of the reciprocal space or, alternatively, between the electron states corresponding to different $\bf{k}$ points. Šmejkal, Sinova, and Jungwirth Smejkal2002 ; Smejkal2002A suggested the term altermagnet for the materials where the magnetic compensation takes place between different $\bf{k}$ points. The absence of the spin degeneracy at the same $\bf{k}$ point can be treated as the spin splitting of the electron states at this point. This property of altermagnets has important physical consequences attracting considerable research attention to this class of materialsSmejkal2002A ; Mazin2002 ; Bai2024 ; Reimers2024 ; Krempasky_MnTe_2024 ; Mazin2023 ; fedchenko2024 .
@@
Among the special properties of altermagets is the chirality splitting of the magnon states Smejkal2023 ; Bai2024 ; Sodequist2024 ; Liu2024 . Each AFM magnon brings either positive or negative magnetization to the system. If in Fig. 1(c) $\theta_{A}$$>$$\theta_{B}$ the magnon gives negative contribution to the magnetization whereas for a magnon with $\theta_{A}$$<$$\theta_{B}$ the contribution to the magnetization is positive. These two types of magnons have opposite chiralities. The two types of magnons are obtained in both Heisenberg model (see, e.g., Ref. Rezende2019, ) and first-principles calculation of the dynamical spin susceptibility (see, e.g., Ref. Sandratskii2012, ). We will refer to these two types of magnons as magnons of type A and B according to the larger of two angles $\theta_{A}$ and $\theta_{B}$. The net magnetization of an AFM at nonzero temperatures remains zero. This property is the consequence of the mutual compensation of the magnons of opposite chiralities which results from their symmetry-determined energy degeneracy. Again there are two possibilities. The chirality compensation can take place either at each magnon wave vector $\bf{q}$ or only between magnons with different wave vectors. The latter situation takes place in altermagnets and may be referred to as chirality splitting of magnons with a given wave vector $\bf{q}$. The nature of the chirality degeneracy of the magnons is very different compared to the spin degeneracy of the electron states discussed above. However, in both cases the two symmetry questions to address are similar: First, which symmetry operations are responsible for the degeneracy of the magnons with opposite chiralities, and, second, does the degeneracy take place between the magnons with the same wave vector or with different wave vectors? As will be discussed, the answers to these questions for magnons are closely related to those for the electron states.
@@
Our choice of altermagnet MnTe as the object of the application of the method has following reasons. First, as a particular case of a two-sublattice AFM containing nonmagnetic atoms (Fig. 2) it is complex enough to demonstrate important features of the method. In the case of MnTe, the altermagnetic properties are the consequences of the presence of the Te atoms. Indeed, in an assumed material with removed Te atoms the altermagnetic spin and chirality splittings are absent. Because of the importance of the Te atoms for the altermagnetism the account for their self-consistent response to the changers in the Mn subsystem is an important feature. In this respect, among the questions to address are, first, whether the Te atoms remain equivalent in the magnon states and, second, what can be said about induced Te moments on the basis of symmetry arguments? The magnon-specific symmetry information about nonmagnetic atoms helps to control and accelerate the convergence of the magnon calculation.
@@
Next question addressed in the paper is how the chirality properties of the magnon states of the system are related to the properties of the electron band structures of these states. The deep connection between two different energy characteristics is exposed.
@@
As seen from the above, the symmetry aspects play in the paper an important role. There are three different parts of the work where the symmetry arguments are essential: (i) the formulation of the method of the direct magnon calculation, (ii) the study of the spin-splitting of the electron states in the GS of an altermagnet, and (iii) the study of the chirality-splitting of the spin waves in an altermagnet. The employment of the SSG allows both solving these tasks and the integration of different parts of the study in one coherent physical picture.
@@ II The method of direct DFT-based magnon calculation
As mentioned in the introduction, in Ref. Niu99, it was shown that for any collinear magnet the magnon energy can be calculated applying Eq. (1). The value of $\Delta m_{z}$ corresponds to the number of magnons with given wave vector and $\Delta E$ is their energy. It is expected that there is an interval of $\Delta m_{z}$ where $\Delta E$ is proportional to $\Delta m_{z}$ and the calculation for any $\Delta m_{z}$ from this interval gives the value of the spin wave energy. The both quantities can be taken per unit cell. To obtain the self-consistent magnon state with a given magnetization ${m}_{\circ}$--$\Delta m_{z}$ the minimization of the density functional must be performed under the constraining condition
@@ III Calculation details
The calculations were performed with the augmented spherical waves (ASW) method Williams1979 ; Eyert2012 . The local density approximation (LDA) to the exchange-correlation functional is usedbarth72 . The account for generalized translation symmetry and external magnetic field have been implemented earlier Uhl1992 ; Miyake2018 . Therefore, only limited adaptation of the code were needed. In some of the calculations we used the LDA+$U$ approach in the flavor of Dudarev et al.Dudarev1998 .
@@
The calculations were performed with three different k meshes in the Brillouin zone (BZ) of the material (Fig. 4): $n$$\times$$n$×$\times n$, $n$=10, 20, and 30. Here $n$ is the number of the intervals in which the primitive vectors of the reciprocal lattice were divided. The results of the calculations have shown that the difference between spin wave energies obtained with $n$=20 and $n$=30 is usually small and most of the calculations reported in the paper were performed with $n$=20.
@@
As mentioned above, a straightforward application of Eq. (1) assumes the existence of an interval of $\Delta m_{z}$ where energy increase $\Delta E$ is proportional to magnetization change $\Delta m_{z}$. Our calculations confirmed that usually such a linear dependence exists up to rather large deviation angles of the Mn atomic moments from the magnetization axis. Figure 5 shows typical characters of the dependences between calculated quantities and demonstrates high stability of the calculated magnon energy with respect to the value of the deviation of the Mn moments from the magnetization axis. The results of the calculations presented in the figure were obtained for wave vector q=(0,0,0.3). Here and in the rest of the paper we give the $x$ and $y$ components of the reciprocal space vectors in units of $\frac{2\pi}{a}$ and the $z$ components in units of $\frac{2\pi}{c}$. We calculated constraining field $h$ stabilizing the deviations of the Mn moments of sublattice A at following values: $\theta_{A}$=5,10,20,30. Figure 5(a) gives calculated deviations of the moments of sublattice B, $\theta_{B}$. Figure (b) shows corresponding increase in energy $\Delta E$. The difference between $\theta_{A}$ and $\theta_{B}$ is the source of the magnetization change $\Delta m_{z}$ shown in figure (c). In figure (d) we present the dependence of $\Delta E$ on $\Delta m_{z}$ which is very close to a linear one. The ratio of these quantities giving the magnon energy is with a good accuracy independent of $\Delta m_{z}$. In most of the calculations presented in the paper we used $\theta_{A}$=20. On the other hand, in the course of LDA+$U$ calculations we dealt with the cases where strong deviation from the simple dependence of $\Delta E$ on the deviation angle of the Mn atomic moments has been obtained showing the capacity of the method to reveal an instability in the studied system connected with the deviation of the atomic moments from the magnetization axis. These results are discussed in Sec. IV.7.

### v2

@@ Abstract
We suggest the first method for direct ab initio calculation of adiabatic magnons in complex collinear magnets. The method is based on the density-functional-theory (DFT) calculation under two different constraints: one constraint governs the change of the magnetization with respect to the ground state, and the other is the symmetry constraint responsible for the value of the magnon wave vector. The advantages of the suggested method with respect to the usual approach of mapping of the electron system on the Heisenberg Hamiltonian of interacting magnetic moments are discussed. The performance of the method is demonstrated by the application to an altermagnet MnTe. The altermagnetism introduced as a concept in 2022 is at present an area of highly intensive research. The characteristic feature of altermagnets is the spin splitting of the electron states in reciprocal k space. Among the discovered properties of the altermagnets is the chirality splitting of the magnons in wave vector q space. We suggest an appoach to the study of the symmetry aspects of magnon chirality splitting. We show that both the chirality splitting of the magnons and the altermagnetic spin splitting of the electron states, though very different in their physical nature, have identical patterns in the corresponding wave vector spaces. Since the altermagnetism of MnTe is the consequence of the presence of the Te atoms, an adequate attention is devoted to the symmetry analysis and calculation results for the Te moments induced in the magnon states. In the calculations each magnon is characterized by its own electron band structure. We investigate the transformation of the electron structure in the transition of the material from the collinear ground state to noncollinear magnon states. We show the connection between the properties of magnon band structures and the chirality properties of magnons. In the investigation of the chirality splitting as well as in both the formulation and the application of our method an important role play the aspects of generalized symmetry based on the application of the spin-space groups. The symmetry framework connects in one coherent picture different parts of the consideration: (i) the generalized translational symmetry of the magnons as a crucial condition for their efficient ab-initio calculation, (ii) altermagnetic spin-splitting of the electron states in the ground state, and (iii) chirality splitting and band structures of the magnon excitations.
@@ I Introduction
There are different approaches to the DFT based calculation of magnons. One of the approaches is the calculation of the dynamic magnetic susceptibility in nonuniform transversal magnetic field Savrasov1998 ; Buczek2011 ; Sandratskii2012 ; Odashima2013 ; Lounis2015 ; Skovhus2021 ; Liu2023 . A strong feature of this method is that it allows to investigate not only the structure and the energy of magnons but also their life time resulting from the interaction of magnons with single-electron Stoner excitations. In the case of complex magnetic materials the method is very demanding with respect to both computation techniques and computer resources. A widely used approach to the theoretical study of the adiabatic magnons is the mapping, as an intermediate step, of the electron system on the Heisenberg Hamiltonian of interacting atomic moments. Such a mapping is currently a standard procedure consisting in the DFT based evaluation of the Heisenberg exchange parameters (see, e.g., Refs. force_theorem, ; Mankovsky_parameters2022, ; Turek_parameters2006, ; Lezaic2006, ; Mu2019, ; Martinez2023, ; Sandratskii2007, ; Sodequist2024, ). An efficient method of the mapping was suggested by Liechtenstein et al. force_theorem and is based on the evaluation of the variation of the band energy as the response to the infinitesimal deviation of the atomic moments from the magnetization axis. The possibility to replace the variation of the total energy by the variation of the band energy is based on so-called magnetic force theorem force_theorem . A recent review of the development and applications of the Liechtenstein et al. method is given in Ref. Szilva2023, . In this paper we suggest a direct DFT-based method for the magnon calculation that does not include mapping of the electron system on the Heisenberg Hamiltonian. To our best knowledge this is the first method allowing direct DFT calculations of adiabatic magnons in complex collinear magnets. The method has important features not provided by the standard mapping approach. Among them are the following: First, the fully self-consistent calculation of the magnon structure and energy is performed in contrast to the mapping approach based on the evaluation of the band energy variation from non self-consistent calculation. Second, in the magnon states of complex structures the atoms that are equivalent in the ground state may become inequivalent. The suggested method takes this into account in a consequent self-consistent manner. Third, in the compounds the contribution of the nominally nonmagnetic atoms to the magnon states is self-consistently taken into account. Forth, the method allows to estimate the dependence of the magnon energy on the number of magnons by varying the tilting angles of the atomic moments from the magnetization axis. The approach suggested by Liechtenstein et al. considers an infinitesimal deviation of the moments from the magnetization axis. Fifth, because of the exact treatment of the generalized periodicity of the helical magnetic structures with arbitrary wave vectors our method accounts for the exchange interactions between atoms at arbitrary large distances in contrast to the mapping calculations that often take into account only a few nearest neighbor interactions.
@@
derived in Ref. Niu99, for the magnon energy in an arbitrary collinear magnetcomment_Niu . Here $\Delta E$ is the increase of the energy of the magnon state with respect to the GS and $\Delta m_{z}$ is the magnetization change with respect to the GS magnetization. Equation (1) correlates with the property that one magnon changes the magnetization of the system by 1 $\mu_{B}$. The first constraint imposes the condition on the magnetization of the system by means of introduction of an effective magnetic field.
The second constraint specifies the value of the magnon wave vector. This is a symmetry type of constraint Sandratskii2001 where the invariance of the initial Hamiltonian with respect to the symmetry operation responsible for the desired property reproduces itself during iterations. The symmetry constraint does not need a constraining field. It is important to emphasize that to constrain the magnon wave vector the generalized translational periodicity described by the machinery of spin space groups (SSG) must be imposed. The combination of two different constraints is the characteristic feature of the method.
@@
In the paper, we formulate the method in the form valid for the study of the magnons in any collinear magnet without spin-orbit coupling (SOC). The performance of the method is demonstrated by the application to altermagnets focusing on MnTe as a representative of this class of materials. As the notion of an altermagnet is rather new it is worth to introduce it briefly. The ground state of a two sublattice AFM is characterized by two mutually compensating ferromagnetic sublattices. In terms of the electron structure, the zero net magnetization of an AFM is the result of the mutual compensation of the spin-up and spin-down electron states which is the consequence of the symmetry properties of the material. Importantly, the realization of this spin compensation can be different for different AFM materials: It can take place either at each wave vector $\bf{k}$ of the reciprocal space or, alternatively, between the electron states corresponding to different $\bf{k}$ points. Šmejkal, Sinova, and Jungwirth Smejkal2002 ; Smejkal2002A suggested the term altermagnet for the materials where the magnetic compensation takes place between different $\bf{k}$ points. The absence of the spin degeneracy at the same $\bf{k}$ point can be treated as the spin splitting of the electron states at this point. This property of altermagnets has important physical consequences attracting enormous research attention to this class of materialsSmejkal2002A ; Mazin2002 ; Bai2024 ; Reimers2024 ; Krempasky_MnTe_2024 ; Mazin2023 ; fedchenko2024 .
Among the special properties of altermagets is the chirality splitting of the magnon states Smejkal2023 ; Bai2024 ; Sodequist2024 ; Liu2024 ; Alaei2025 . In Ref. Smejkal2023, the presence of the magnon chirality splitting is connected with the properties of the Heisenberg exchange parameters. A general approach to the chirality splitting applicable in the DFT-based direct calculations has not yet been suggested. Also the discussion of the symmetry pattern of the magnon chirality splitting in the reciprocal wave-vector space and its relation to the symmetry partern of the exchange splitting in the ground-state band structure is not available. To give the answers to these questions is one of the aims of this paper.
@@
Each AFM magnon brings either positive or negative magnetization to the system. If in Fig. 1(c) $\theta_{A}$$>$$\theta_{B}$ the magnon gives negative contribution to the magnetization whereas for a magnon with $\theta_{A}$$<$$\theta_{B}$ the contribution to the magnetization is positive. These two types of magnons have opposite chiralities. The two types of magnons are obtained in both Heisenberg model (see, e.g., Ref. Rezende2019, ) and first-principles calculation of the dynamical spin susceptibility (see, e.g., Ref. Sandratskii2012, ). We will refer to these two types of magnons as magnons of type A and B according to the larger of two angles $\theta_{A}$ and $\theta_{B}$. The net magnetization of an AFM at nonzero temperatures remains zero. This property is the consequence of the mutual compensation of the magnons of opposite chiralities which results from their symmetry-determined energy degeneracy. Again there are two possibilities. The chirality compensation can take place either at each magnon wave vector $\bf{q}$ or only between magnons with different wave vectors. The latter situation takes place in altermagnets and may be referred to as chirality splitting of magnons with a given wave vector $\bf{q}$. The nature of the chirality degeneracy of the magnons is very different compared to the spin degeneracy of the electron states discussed above. However, in both cases the two symmetry questions to address are similar: First, which symmetry operations are responsible for the degeneracy of the magnons with opposite chiralities, and, second, does the degeneracy take place between the magnons with the same wave vector or with different wave vectors? As will be demonstrated, the answers to these questions for magnons are closely related to those for the electron states.
@@
Our choice of altermagnet MnTe as the object of the application of the method has following reasons. First, as a particular case of a two-sublattice AFM containing nonmagnetic atoms (Fig. 2) it is complex enough to demonstrate important features of the method. In the case of MnTe, the altermagnetic properties are the consequences of the presence of the Te atoms. Indeed, in an assumed material with removed Te atoms the altermagnetic spin and chirality splittings are absent. Due to the crucial role of Te atoms in altermagnetism, accounting for their self- consistent response to changes in the Mn subsystem is a key aspect. In this context, key questions to address include the questions of whether the Te atoms remain equivalent in the magnon states and of what symmetry arguments can reveal about the induced Te moments. The magnon-specific symmetry information about nonmagnetic atoms, besides its general physical importance, helps to control and accelerate the convergence of the magnon calculation. These ’technical’ advantages from the symmetry analysis are discussed in Sec. IV.4.
@@
Next question addressed in the paper is how the chirality properties of the magnon states of the system are related to the properties of the electron band structures of the magnon states. To our best knowledge this is the first study of this type. It is based on the possibility to calculate the band structure of the magnon states opened by our method. The deep connection between two different energy characteristics of altermagnets, magnon energies and magnon electron band structures, is exposed.
@@
As seen from the above, the symmetry aspects play in the paper an important role. There are three different parts of the work where the symmetry arguments are essential: (i) the formulation of the method of the direct magnon calculation, (ii) the study of the spin-splitting of the electron states in the GS of an altermagnet, and (iii) the study of the chirality-splitting of the spin waves in an altermagnet. The employment of the SSG allows both solving these tasks and the integration of different parts of the study in one coherent physical picture. We consider this coherent picture uniting very different sides of the consideration as one of important results of the paper. As discussed below, in part (ii) the employment of usual space groups can be technically sufficient. In other parts and in the formation of a general picture the use of the generalized spin-space groups is essential.
@@ II The method of direct DFT-based magnon calculation
As mentioned in the introduction, in Ref. Niu99, it was shown that for any collinear magnet the magnon energy can be presented within the framework of the DFT theory in the form given by Eq. (1) (see also Ref. comment_Niu, ). The value of $\Delta m_{z}$ corresponds to the number of magnons with given wave vector and $\Delta E$ is their energy. It is expected that there is an interval of $\Delta m_{z}$ where $\Delta E$ is proportional to $\Delta m_{z}$ and the calculation for any $\Delta m_{z}$ from this interval gives the value of the spin wave energy. The both quantities can be considered per unit cell. To obtain the self-consistent magnon state with a given magnetization ${m}_{\circ}$--$\Delta m_{z}$ the minimization of the density functional must be performed under the constraining condition
@@ III Calculation details
The calculations were performed with the augmented spherical waves (ASW) method Williams1979 ; Eyert2012 . The local density approximation (LDA) to the exchange-correlation functional is usedbarth72 . The account for generalized translation symmetry and external magnetic field have been implemented earlier Uhl1992 ; Miyake2018 . Therefore, only limited adaptation of the code were needed.
@@
The calculations were performed with three different k meshes in the Brillouin zone (BZ) of the material (Fig. 4): $n$$\times$$n$×$\times n$, $n$=10, 20, and 30. Here $n$ is the number of the intervals in which the primitive vectors of the reciprocal lattice were divided. The results of the calculations have shown that the difference between spin wave energies obtained with $n$=20 and $n$=30 is usually small and most of the calculations reported in the paper were performed with $n$=20. In the calculations we used experimental lattice parametersSzuszkiewicz2006 $a=$ 4.15 Å, $c=$ 6.71 Å.
@@
As mentioned above, a straightforward application of Eq. (1) assumes the existence of an interval of $\Delta m_{z}$ where energy increase $\Delta E$ is proportional to magnetization change $\Delta m_{z}$. Our calculations confirmed that usually such a linear dependence exists up to rather large deviation angles of the Mn atomic moments from the magnetization axis. Figure 5 shows typical characters of the dependences between calculated quantities and demonstrates high stability of the calculated magnon energy with respect to the value of the deviation of the Mn moments from the magnetization axis. The results of the calculations presented in the figure were obtained for wave vector q=(0,0,0.3). Here and in the rest of the paper we give the $x$ and $y$ components of the reciprocal space vectors in units of $\frac{2\pi}{a}$ and the $z$ components in units of $\frac{2\pi}{c}$. We calculated constraining field $h$ stabilizing the deviations of the Mn moments of sublattice A at following values: $\theta_{A}$=5,10,20,30. Figure 5(a) gives calculated deviations of the moments of sublattice B, $\theta_{B}$. Figure (b) shows corresponding increase in energy $\Delta E$. The difference between $\theta_{A}$ and $\theta_{B}$ is the source of the magnetization change $\Delta m_{z}$ shown in figure (c). In figure (d) we present the dependence of $\Delta E$ on $\Delta m_{z}$ which is very close to a linear one. The ratio of these quantities giving the magnon energy is with a good accuracy independent of $\Delta m_{z}$. In most of the calculations presented in the paper we used $\theta_{A}$=20. On the other hand, performing model calculations aimed at testing our method we obtained also the cases where strong deviation from the simple dependence of $\Delta E$ on the deviation angle of the Mn atomic moments was registered showing the capacity of the method to reveal an instability connected with the deviation of the atomic moments from the magnetization axis. These results are briefly presented in Sec. IV.7.